# Magnetic field measurements in a sample of Class I and Flat-Spectrum protostars observed with SPIRou

### 18 Mar 2026 v1

@@ 4.4 Comparison with the total magnetic fields
Out of their 32 sources, Flores et al. (2024) was able to measure non-null ¡BB¿ in 27 of them, and report upper limits for the 5 remaining sources between 0.8 and 1.2 kG. Among the 27 sources with a measured magnetic field, they report an average magnetic field strength of approximately 2 kG, consistent with the values typically found in more evolved Class II sources (Johns-Krull, 2007). These measurements, based on Zeeman broadening, are sensitive to the total magnetic field, including both large and small-scale components. In contrast, our spectropolarimetric method is sensitive only to large-scale magnetic fields. These two approaches probe different components. Indeed, while Zeeman broadening captures the cumulative surface magnetic field regardless of orientation, spectropolarimetry detects only the longitudinal component of the field. This means that regions of opposite polarity on the stellar surface can cancel each other out. If the stellar magnetic field is only organised on small scales, or is a combination of large- and small-scales fields, therefore the spectropolarimetric method leads to an underestimation of the total magnetic field (See et al., 2019; Lehmann et al., 2019).
As already mentioned, our snapshot sample cannot be used to determine the topology and strength of the large scale magnetic field, and therefore estimate its contribution to the total magnetic field measured through Zeeman broadening. However, if we assume that our magnetic sample contains only dipolar fields, then we can use the statistical relationship between the cumulative longitudinal magnetic field ($B_{\ell}$¿rms) of a magnetic sample, and its average total magnetic strength ¡BB¿av. Kochukhov (2024) determined a statistical relationship between the two parameters by running many simulations of dipoles at the surface of an A0 main-sequence star for random magnetic obliquities and inclinations, and of a linear limb-darkening parameter of 0.5. The average value of limb-darkening parameter of our sample is 0.4 (Claret and Bloemen, 2011). When correcting for our different limb-darkening value, the relationship can be written as:
$\langle B\rangle_{\mathrm{av}}=4.04^{+1.1}_{-0.5}\,\langle B_{\ell}\rangle_{\rm rms}$ (2)
@@
Figure 5: Projected rotational velocities $v\sin i$, as a function of effective temperatures measured in this work (blue), and in the work of Flores et al. (2024) (red). The connected points with the dashed-line indicate the common stars in both samples.
@@ 5 Conclusions
We have carried out a detailed analysis of 14 Class I and FS protostars using SPIRou, aiming at characterising their stellar and large-scale magnetic properties. When including V347 Aur, the final statistical sample comprises 15 objects. When possible, we derived effective temperatures, and using LSD profiles, we measured the detection probability of a Zeeman signature in the Stokes $V$ profiles, the longitudinal magnetic fields, the projected rotational velocities, and the radial velocities. Our results can be summarized as follows:
@@
    We detect large-scale magnetic fields in 6 out of 15 protostars, with absolute longitudinal field strengths ranging from approximately 80 G to 200 G. In contrast, nine objects show no clear magnetic signature.
@@
    Effective temperatures for our sample of Class I and FS sources range from 3250 K to 5000 K, with a median value of approximately 4000 K.
@@
    We derive upper limits on the dipolar field strength, with a median value below 3 kG, indicating that strong magnetic fields may still be present but are below our detection threshold when no magnetic field is detected.
@@
    The longitudinal magnetic field values of our magnetic sample do not seem to reach the higher levels observed in their evolutionary descendants, i.e. the fully convective CTTS.
@@
Importantly, recent studies have pushed the observational frontier to even earlier stages. Although Class 0 protostars are even more deeply embedded and challenging to observe, recent studies have begun to detect photospheric absorption lines in these objects (Greene et al., 2018; Le Gouellec et al., 2025). These detections open promising prospects for future spectropolarimetric studies of Class 0 sources, offering the potential to provide critical constraints on the initial conditions of dynamo generation and accretion physics, and trace magnetic field evolution from the very first stages of stellar formation. Class 0 objects have similar accretion diagnostics as in the more evolved Class I and II protostars (Laos et al., 2021; Le Gouellec et al., 2024), maybe suggesting that strong and organised stellar magnetic fields are present.
This study demonstrates the capabilities of SPIRou to probe the large-scale magnetic fields of embedded protostars, making it a powerful tool to characterise magnetism in the early stages of star formation. The next step is to perform ZDI on our magnetic sample. This will permit to unveil the topology of the magnetic field in Class I and FS objects, which has never been done to this date. These topologies will be a key element in understanding the evolution of magnetic fields in protostars and of the star-disk interactions.
@@ Acknowledgements.
We want to thank the referee for helpful comments which significantly improved the clarity of the manuscript. This work has received funding from the French Agence Nationale de la Recherche (ANR) through the project PROMETHEE (ANR-22-CE31-0020). This research has made use of the SIMBAD database, the VizieR catalogue access tool, and the Aladin sky atlas, CDS, Strasbourg Astronomical Observatory, France. OK acknowledges support by the Swedish Research Council (grant agreement no. 2023-03667) and the Swedish National Space Agency. AK acknowledges support by the NKFIH NKKP grant ADVANCED 149943 and the NKFIH excellence grant TKP2021-NKTA-64. Project no.149943 has been implemented with the support provided by the Ministry of Culture and Innovation of Hungary from the National Research, Development and Innovation Fund, financed under the NKKP ADVANCED funding scheme. CPF acknowledges funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No. 101079231 (EXOHOST) and from UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee (grant number 10051045).
@@ Appendix A The dipole strength upper limit method
The Nelder-Mead algorithm is particularly well-suited for minimising functions that are difficult to differentiate, such as a $\chi^{2}$ function. To better constrain the initial parameters, we estimated them with a Gaussian fit, which provides a reliable estimate of the most probable initial values. To estimate the uncertainties associated with the optimised parameters, we performed a detailed $\chi^{2}$ analysis. Starting from the optimal parameter values, each parameter was varied individually by 20% around its best-fit value while keeping the other parameters fixed. For each parameter, the $\chi^{2}$ function was evaluated at 20 evenly spaced points within this range. The minimum total $\chi^{2}$ value is computed, the upper and lower bounds of the parameter were found by locating the points where $\chi^{2}=\chi^{2}_{\text{min}}+1$. These bounds define the asymmetric uncertainties ($\sigma_{-}$ and $\sigma_{+}$) for each parameter, reflecting the sensitivity of the $\chi^{2}$ function to changes in the parameter values. To further validate the reliability of the uncertainties, we compared the results with other methods, such as Monte Carlo simulations and Probability Density Function analysis. Additionally to the one-dimensional analysis, we performed a two-dimensional $\chi^{2}$ analysis to examine correlations between pairs of parameters. In this case, the criterion for the 68% confidence region corresponds to $\Delta\chi^{2}=2.4$, as there are two free parameters. These complementary approaches confirmed the robustness of the uncertainty estimates, ensuring that the reported confidence intervals are both accurate and reliable.
Figure 6: Example of dipolar magnetic detection probability curve as a function of the dipole strength. The colour curves represent the probability for each individual observation, while the black curve represents the combined probability for all observations. The 90 percent threshold is used to determine the field strength upper limit.
@@
Next, a Monte-Carlo method is used to simulate a grid of 1000 synthetic $V$ profiles for each observation, where random values are assigned to the model parameters. Each model uses a random inclination angle $i$, obliquity angle $\beta$, and rotational phase $\Omega$, with various values of the polar magnetic fields $B_{\text{pol}}$ using the same mean Landé factor (1.2) and wavelength (1660 nm). The local line profiles are computed using an instrumental broadening corresponding to a resolving power of 70 000 and a macroturbulent velocity of 5 km s-1. Indeed, the shapes of the wings of the LSD $I$ profiles are not compatible with a null macroturbulent velocity. Furthermore, if we let it be free in the fit, we obtain values that are unphysical. We, therefore, made a compromise and chose to fix it to a value that reproduces well the wings of the LSD profiles, but is not too high. We should caution that the computation of the LSD profiles themselves may produce artificially broadened wings because of, e.g. line blending, or depressed continuum. It, therefore, can be higher than the value needed to fit the individual spectral lines for the temperature determination in Sect. 3.1. This value should, therefore, be interpreted with caution. The surface integration of the local profiles is done using a linear limb-darkening coefficient. These values have been extracted from the Claret and Bloemen (2011) database for the corresponding effective temperatures of our stars (Sect. 3.1), assuming a surface gravity of 3.5, and solar abundances. Because our LSD masks are dominated by H-band lines, we choose the coefficients computed in this band, and we used the mean values of both methods (Least-Squares or Flux Conservation) used by Claret and Bloemen (2011) to compute them.
@@
Gaussian noise is added to each profile. This process is repeated for all the observations, with the 1000 models of VV profiles enabling the calculation of the detection probability for a range of magnetic field strengths. For each set of models, the probability of detecting a magnetic field in each simulated $V$ profile is computed. This allows us to determine whether a magnetic field is detected or not, using the False Alarm Probability. We ensured that the FAP does not exceed a prescribed value, using FAP = $10^{-3}$ for marginal detection (Donati et al. 1997). We chose the marginal criterion to be consistent with previous publications, and be able to compare with other samples. We also compute the upper limit using definite criterion. However, the choice of this criterion is arbitrary, as all values will be affected in a similar way. For each value of the field strength, we calculated the number of detections across several models for each observation and plotted the detection rates as a function of field strength. A detection rate of 90 percent is required to statistically consider that the field would be detected in one observation. This enables us to determine the upper limit of the field strength, below which we are unable to detect it in our observations. Assuming the observations are all independent, we can compute the combined probability following the method outlined by Neiner et al. (2015), for various polar field strengths, and apply the 90 percent detection threshold to derive the upper limit. By analysing this curve, we can determine a value of $B_{\text{lim}}$. This corresponds either to a profile containing only noise or one in which the signature of a weak magnetic field is hidden within the spectral noise in Stokes V (see Fig 6).
@@ Appendix B Fitting example with ZeeTurbo
Figure 7: ZeeTurbo fitting process for the first observation of IRAS 04113+2758 N. The observed spectrum is shown in black, and the best fit model in red. Blue boxes indicate the regions used for parameter fitting, and the corresponding absorption lines are labelled.
@@ Appendix C The effect of veiling on the uncertainties of $B_{\ell}$
As our sources can be strongly affected by veiling, i.e. a continuum excess that adds to the stellar continuum, we explore its effect on the uncertainties on the $B_{\ell}$ measurement. We can re-write Eq.(1) as:
@@
and $\alpha=-2.14\times 10^{11}/(\lambda zc)\approx-358$ G km-1 s when using the normalisation values of $\lambda$ and $z$ of our LSD profiles. The numerical computation of $X$ and $Y$, and of their uncertainties $\sigma_{X}$ and $\sigma_{Y}$ can be written:
@@
where $\sigma_{I}$ and $\sigma_{V}$ are the uncertainties on the unnormalised $I$ and $V$ spectra, and $N$ is the number of pixels within the photospheric profile. The uncertainty on $B_{\ell}$ is therefore:
@@
Let’s consider a star that we observe twice. During the first observation the star is not affected by veiling, and emits a spectrum $I_{1}(v)=I_{0}(v)$ with a continuum $I_{\rm 1c}=I_{\star}$. During the second observation the star is contaminated with an additional excess continuum $I_{\rm ex}(v)=rI_{\star}(v)$, where $r$ is defined as the veiling parameter. We will assume that $r$ does not depend on the wavelength, and that the region emitting the veiling does not emit circular polarisation. During the second observations we therefore record the spectrum $I_{2}(v)=I_{0}(v)+I_{\rm ex}(v)$, with a continuum $I_{\rm 2c}=I_{\star}+I_{\rm ex}$. Both polarised spectra are assumed identical, and an associated polarised spectrum $V_{2}(v)=V_{1}(v)=V(v)$. $X$ and $Y$ for both observations are therefore:
@@
As the continuum in the LSD profiles is not dependent on $v$, when dividing $X$ by $Y$ in both observations, the numerators cancel, and $X_{1}/Y_{1}=X_{2}/Y_{2}$, hence:
$\hskip 10.00002ptB_{\ell 1}=B_{\ell 2}=B_{\ell}$
To explore how veiling affects the computation of the uncertainties, let’s assume first that both observations were obtained in identical weather conditions, with the same exposure time. The noise is therefore identical in both observations: $\sigma_{I1}=\sigma_{I2}=\sigma_{I}$, and $\sigma_{V1}=\sigma_{V2}=\sigma_{V}$
@@
When dividing $\sigma_{X}$ by $X$ and $\sigma_{Y}$ by $Y$, the denominators of $X$ and $Y$ cancel, and therefore:
$\hskip 10.00002pt\frac{\sigma_{B1}}{B_{\ell 1}}=\frac{\sigma_{B2}}{B_{\ell 2}}=\frac{\sigma_{B}}{B_{\ell}}$
@@
Let’s now assume a second case where we observe two identical stars of the same magnitude. One is veiled, and the other one is not. However, before scheduling the observation, we do not know that the second star is veiled. We therefore observe them with the same exposure time, in the same weather conditions. Because the magnitudes are the same, the recorded fluxes for both stars are the same, and therefore $(S/N)_{2}=(S/N)_{1}$. The uncertainties on the flux from the stars, however, are not the same:
@@
Substituting these expressions into the definitions of $\sigma_{X}$ and $\sigma_{Y}$ give:
@@
Hence, for the veiled observation we obtain:

### 22 Mar 2026 v2

@@ 4.4 Comparison with the total magnetic fields
Out of their 32 sources, Flores et al. (2024) was able to measure non-null ¡BB¿ in 27 of them, and report upper limits for the 5 remaining sources between 0.8 and 1.2 kG. Among the 27 sources with a measured magnetic field, they report an average magnetic field strength of approximately 2 kG, consistent with the values typically found in more evolved Class II sources (Johns-Krull, 2007). These measurements, based on Zeeman broadening, are sensitive to the total magnetic field, including both large and small-scale components. In contrast, our spectropolarimetric method is sensitive only to large-scale magnetic fields. These two approaches probe different components. While Zeeman broadening captures the cumulative surface magnetic field regardless of orientation, spectropolarimetry detects only the longitudinal component of the field. This means that regions of opposite polarity on the stellar surface can cancel each other out. If the stellar magnetic field is only organised on small scales, or is a combination of large- and small-scales fields, the spectropolarimetric method leads to an underestimation of the total magnetic field (See et al., 2019; Lehmann et al., 2019).
As already mentioned, our snapshot sample cannot be used to determine the topology and strength of the large scale magnetic field, and therefore estimate its contribution to the total magnetic field measured through Zeeman broadening. However, if we assume that our magnetic sample contains only dipolar fields, then we can use the statistical relationship between the cumulative longitudinal magnetic field ($B_{\ell}$¿rms) of a magnetic sample, and its average total magnetic strength ¡BB¿av. Kochukhov (2024) determined a statistical relationship between the two parameters by running many simulations of dipoles at the surface of an A0 main-sequence star for random magnetic obliquities and inclinations, and of a linear limb-darkening parameter of 0.5. The average value of limb-darkening parameter of our sample is 0.4 (Claret and Bloemen, 2011). When correcting for our different limb-darkening value, the relationship can be written as
$\langle B\rangle_{\mathrm{av}}=4.04^{+1.1}_{-0.5}\,\langle B_{\ell}\rangle_{\rm rms},$ (2)
@@
Figure 5: Projected rotational velocities $v\sin i$as a function of effective temperatures measured in this work (blue) and in the work of Flores et al. (2024, red). The points connected by the dashed line indicate the common stars in the two samples.
@@ 5 Conclusions
We have carried out a detailed analysis of 14 Class I and FS protostars using SPIRou, aiming at characterising their stellar and large-scale magnetic properties. When including V347 Aur, the final statistical sample comprises 15 objects. When possible, we derived effective temperatures, and using LSD profiles, we measured the detection probability of a Zeeman signature in the Stokes $V$ profiles, the longitudinal magnetic fields, the projected rotational velocities, and the radial velocities. Our results can be summarised as follows:
@@
    We detected large-scale magnetic fields in 6 out of 15 protostars, with absolute longitudinal field strengths ranging from approximately 80 G to 200 G. In contrast, nine objects show no clear magnetic signature.
@@
    The effective temperatures for our sample of Class I and FS sources range from 3250 K to 5000 K, with a median value of approximately 4000 K.
@@
    We derived upper limits on the dipolar field strength, with a median value below 3 kG, indicating that strong magnetic fields may still be present, but are below our detection threshold when no magnetic field is detected.
@@
    The longitudinal magnetic field values of our magnetic sample do not seem to reach the higher levels observed in their evolutionary descendants, i.e. the fully convective CTTSs.
@@
Importantly, recent studies have pushed the observational frontier to even earlier stages. Although Class 0 protostars are even more deeply embedded and challenging to observe, recent studies have begun to detect photospheric absorption lines in these objects (Greene et al., 2018; Le Gouellec et al., 2025). These detections open promising prospects for future spectropolarimetric studies of Class 0 sources, offering the potential to provide critical constraints on the initial conditions of dynamo generation and accretion physics, and trace magnetic field evolution from the very first stages of stellar formation. Class 0 objects have accretion diagnostics that are similar to those in the more evolved Class I and II protostars (Laos et al., 2021; Le Gouellec et al., 2024), perhaps suggesting that strong and organised stellar magnetic fields are present.
This study demonstrates the capabilities of SPIRou to probe the large-scale magnetic fields of embedded protostars, making it a powerful tool to characterise magnetism in the early stages of star formation. The next step is to perform ZDI on our magnetic sample. This will unveil the topology of the magnetic field in Class I and FS objects, which has not been done to date. These topologies will be a key element in understanding the evolution of magnetic fields in protostars and of the star–disk interactions.
@@ Acknowledgements.
We want to thank the referee for helpful comments which significantly improved the clarity of the manuscript. This work has received funding from the French Agence Nationale de la Recherche (ANR) through the project PROMETHEE (ANR-22-CE31-0020). This research has made use of the SIMBAD database, the VizieR catalogue access tool, and the Aladin sky atlas, CDS, Strasbourg Astronomical Observatory, France. OK acknowledges support by the Swedish Research Council (grant agreement no. 2023-03667) and the Swedish National Space Agency. AK acknowledges support by the NKFIH NKKP grant ADVANCED 149943 and the NKFIH excellence grant TKP2021-NKTA-64. Project no.149943 has been implemented with the support provided by the Ministry of Culture and Innovation of Hungary from the National Research, Development and Innovation Fund, financed under the NKKP ADVANCED funding scheme. CPF acknowledges funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No. 101079231 (EXOHOST) and from UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee (grant number 10051045). H.N. acknowledges funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 101019653). V.J.M.L.G. acknowledges support by the Spanish program Unidad de Excelencia María de Maeztu CEX2020-001058-M, financed by MCIN/AEI/10.13039/501100011033, and by the MaX-CSIC Excellence Award MaX4-SOMMA-ICE. V.J.M.L.G. acknowledges support by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 101098309 - PEBBLES).
@@ Appendix A The dipole strength upper limit method
The Nelder-Mead algorithm is particularly well suited for minimising functions that are difficult to differentiate, such as a $\chi^{2}$ function. To better constrain the initial parameters, we estimated them with a Gaussian fit, which provides a reliable estimate of the most probable initial values. To estimate the uncertainties associated with the optimised parameters, we performed a detailed $\chi^{2}$ analysis. Starting from the optimal parameter values, each parameter was varied individually by 20% around its best-fit value while keeping the other parameters fixed. For each parameter, the $\chi^{2}$ function was evaluated at 20 evenly spaced points within this range. The minimum total $\chi^{2}$ value is computed, the upper and lower bounds of the parameter were found by locating the points where $\chi^{2}=\chi^{2}_{\text{min}}+1$. These bounds define the asymmetric uncertainties ($\sigma_{-}$ and $\sigma_{+}$) for each parameter, reflecting the sensitivity of the $\chi^{2}$ function to changes in the parameter values. To further validate the reliability of the uncertainties, we compared the results with other methods, such as Monte Carlo simulations and Probability Density Function analysis. Additionally to the one-dimensional analysis, we performed a two-dimensional $\chi^{2}$ analysis to examine correlations between pairs of parameters. In this case, the criterion for the 68% confidence region corresponds to $\Delta\chi^{2}=2.4$, as there are two free parameters. These complementary approaches confirmed the robustness of the uncertainty estimates, ensuring that the reported confidence intervals are both accurate and reliable.
Figure 6: Example of dipolar magnetic detection probability curve as a function of the dipole strength. The coloured curves represent the probability for each individual observation, while the black curve represents the combined probability for all observations. The 90 per cent threshold is used to determine the field strength upper limit.
@@
Next, a Monte-Carlo method is used to simulate a grid of 1000 synthetic $V$ profiles for each observation, where random values are assigned to the model parameters. Each model uses a random inclination angle $i$, obliquity angle $\beta$, and rotational phase $\Omega$, with various values of the polar magnetic fields $B_{\text{pol}}$ using the same mean Landé factor (1.2) and wavelength (1660 nm). The local line profiles are computed using an instrumental broadening corresponding to a resolving power of 70 000 and a macroturbulent velocity of 5 km s-1. The shapes of the wings of the LSD $I$ profiles are not compatible with a null macroturbulent velocity. Furthermore, if we let it be free in the fit, we obtain values that are unphysical. We therefore made a compromise and chose to fix it to a value that reproduces well the wings of the LSD profiles, but is not too high. We should caution that the computation of the LSD profiles themselves may produce artificially broadened wings, for example because of line blending, or may produce depressed continuum. It can therefore be higher than the value needed to fit the individual spectral lines for the temperature determination in Sect. 3.1. This value should therefore be interpreted with caution. The surface integration of the local profiles is done using a linear limb-darkening coefficient. These values have been extracted from the Claret and Bloemen (2011) database for the corresponding effective temperatures of our stars (Sect. 3.1), assuming a surface gravity of 3.5, and solar abundances. Because our LSD masks are dominated by H-band lines, we choose the coefficients computed in this band, and we used the mean values of both methods (least-squares or flux conservation) used by Claret and Bloemen (2011) to compute them.
@@
Gaussian noise is added to each profile. This process is repeated for all the observations, with the 1000 models of VV profiles enabling the calculation of the detection probability for a range of magnetic field strengths. For each set of models, the probability of detecting a magnetic field in each simulated $V$ profile is computed. This allows us to determine whether a magnetic field is detected or not, using the FAP. We ensured that the FAP does not exceed a prescribed value, using FAP = $10^{-3}$ for marginal detection (Donati et al. 1997). We chose the marginal criterion to be consistent with previous publications, and be able to compare with other samples. We also compute the upper limit using definite criterion. However, the choice of this criterion is arbitrary, as all values will be affected in a similar way. For each value of the field strength, we calculated the number of detections across several models for each observation and plotted the detection rates as a function of field strength. A detection rate of 90 per cent is required to statistically consider that the field would be detected in one observation. This enables us to determine the upper limit of the field strength, below which we are unable to detect it in our observations. Assuming the observations are all independent, we can compute the combined probability following the method outlined by Neiner et al. (2015), for various polar field strengths, and apply the 90 per cent detection threshold to derive the upper limit. By analysing this curve, we can determine a value of $B_{\text{lim}}$. This corresponds either to a profile containing only noise or to one in which the signature of a weak magnetic field is hidden within the spectral noise in Stokes V (see Fig 6).
@@ Appendix B Fitting example with ZeeTurbo
Figure 7: ZeeTurbo fitting process for the first observation of IRAS 04113+2758 N. The observed spectrum is shown in black, and the best-fit model in red. The blue shading indicates the regions used for parameter fitting, and the corresponding absorption lines are labelled.
@@ Appendix C The effect of veiling on the uncertainties of $B_{\ell}$
As our sources can be strongly affected by veiling, i.e. a continuum excess that adds to the stellar continuum, we explore its effect on the uncertainties on the $B_{\ell}$ measurement. We can re-write Eq. (1) as
@@
and $\alpha=-2.14\times 10^{11}/(\lambda zc)\approx-358$ G km-1 s when using the normalisation values of $\lambda$ and $z$ of our LSD profiles. The numerical computation of $X$ and $Y$, and of their uncertainties $\sigma_{X}$ and $\sigma_{Y}$ can be written
@@
where $\sigma_{I}$ and $\sigma_{V}$ are the uncertainties on the unnormalised $I$ and $V$ spectra, and $N$ is the number of pixels within the photospheric profile. The uncertainty on $B_{\ell}$ is therefore
@@
We consider a star that we observe twice. During the first observation the star is not affected by veiling, and emits a spectrum $I_{1}(v)=I_{0}(v)$ with a continuum $I_{\rm 1c}=I_{\star}$. During the second observation the star is contaminated with an additional excess continuum $I_{\rm ex}(v)=rI_{\star}(v)$, where $r$ is defined as the veiling parameter. We assume that $r$ does not depend on the wavelength, and that the region emitting the veiling does not emit circular polarisation. During the second observations we therefore record the spectrum $I_{2}(v)=I_{0}(v)+I_{\rm ex}(v)$, with a continuum $I_{\rm 2c}=I_{\star}+I_{\rm ex}$. Both polarised spectra are assumed identical, and an associated polarised spectrum $V_{2}(v)=V_{1}(v)=V(v)$. $X$ and $Y$ for both observations are therefore
@@
As the continuum in the LSD profiles is not dependent on $v$, when dividing $X$ by $Y$ in both observations, the numerators cancel, and $X_{1}/Y_{1}=X_{2}/Y_{2}$, hence
$\hskip 10.00002ptB_{\ell 1}=B_{\ell 2}=B_{\ell}.$
To explore how veiling affects the computation of the uncertainties, we assume first that both observations were obtained in identical weather conditions, with the same exposure time. The noise is therefore identical in both observations: $\sigma_{I1}=\sigma_{I2}=\sigma_{I}$, and $\sigma_{V1}=\sigma_{V2}=\sigma_{V}$
@@
When dividing $\sigma_{X}$ by $X$ and $\sigma_{Y}$ by $Y$, the denominators of $X$ and $Y$ cancel, and therefore
$\hskip 10.00002pt\frac{\sigma_{B1}}{B_{\ell 1}}=\frac{\sigma_{B2}}{B_{\ell 2}}=\frac{\sigma_{B}}{B_{\ell}}.$
@@
We now assume a second case where we observe two identical stars of the same magnitude. One is veiled, and the other one is not. However, before scheduling the observation, we do not know that the second star is veiled. We therefore observe them with the same exposure time, in the same weather conditions. Because the magnitudes are the same, the recorded fluxes for both stars are the same, and therefore $(S/N)_{2}=(S/N)_{1}$. The uncertainties on the flux from the stars, however, are not the same:
@@
Substituting these expressions into the definitions of $\sigma_{X}$ and $\sigma_{Y}$ give
@@
Hence, for the veiled observation we obtain